A simple random sample of 5 months of sales data provided the following information 

Month 1 2 3 4 5 

Units Sold 96 110 82 96 91 


a. Develop a point estimate of the population mean number of units sold per month. 

b. Develop a point estimate of the population standard deviation (to 2 decimals).
Answer #1

x : means the unit sold

n = total number of month = 5

point estimate of the population mean = sample mean

sample \ mean = \bar X = \frac{\sum x} {n}

sample \ mean = \bar X = \frac{96+110+82+96+91} {5}

sample \ mean = \bar X = \frac{475}{5} = 95

sample \ mean = 95

Answer for "a" part

point estimate of the population mean number of unit sold per month

95

Part "b "

The point estimate of the population standard devation = sample standard deviation

sample \ standard \ deviation = s = \sqrt \frac{\sum (x - sample \ mean)^2 } {n-1 }

s = \sqrt \frac{ (96- 95)^2 + (110 - 95)^2+ ( 82-95 )^2+ (96-95)^2+(91-95)^2} {5 -1 }

s = \sqrt \frac{ (1)^2 + (-15)^2+ ( -13 )^2+ (1)^2+(-4)^2} {4 }

s = \sqrt \frac{ 1+225+169+1+16} {4 } = \sqrt \frac{ 412 } {4 }

s = \sqrt{103} = 10.14889157

sample \ standard \ deviation = s = 10.14889157

Round the sample standard deviation up to 2 decimal place

sample \ standard \ deviation = s = 10.15

Final answer for b:-

point estimate of the population standard deviation = 10.15